OpenStudy (anonymous):

10/3=k+12 9 what does k eaquel need help will give meadle to the person

OpenStudy (anonymous):

k + 12- 12 = 10/3 - 12 k = 10/3 - 12 k = 10/3 - 36/3 k = -26/3

OpenStudy (anonymous):

thank u